. An elastic band is stretched a distance of 15.0 cm by a force of 10.0 N. a) Calculate the work done in stretching the elastic
b) How much elastic potential energy is stored in the elastic?
c) A 1.00-g object is placed on the elastic, and the elastic is released in a vertical direction. How much kinetic energy is transferred to the object by the elastic?
d) What was the speed of the object at the point that it leaves the elastic?
e) How high will the object in question (d) rise?
1
Expert's answer
2020-12-21T13:40:17-0500

a)


"W=Fx=(10)(0.15)=1.5\ J"

b)


"E=W=1.5\ J"

c)


"K=1.5\ J"

d)


"K=0.5mv^2\\1.5=0.5(0.001)v^2\\v=54.8\frac{m}{s}"

e)


"E=mgh\\1.5=(0.001)(9.8)h\\h=153\ m"